OpenStudy (anonymous):

Simplify: i(sqrt-98) -( sqrt98)

OpenStudy (anonymous):

\[i\sqrt{-98}-\sqrt{98}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2sqrt(98)

OpenStudy (anonymous):

\[\sqrt{-98}=\sqrt{98}i\] giving \[i^2\sqrt{98}-\sqrt{98}=-\sqrt{98}-\sqrt{98}=-2\sqrt{98}\]

OpenStudy (anonymous):

not done yet however!

OpenStudy (anonymous):

11i

OpenStudy (anonymous):

\[98=2\times 49\] so \[\sqrt{98}=7\sqrt{2}\]

OpenStudy (anonymous):

therefore you "final answer is \[-14\sqrt{2}\]

OpenStudy (anonymous):

11i?